The area of a square with side s is A(s)=s². What is the rate of change of the area of a square with respect to its side length when s=8?
1
Expert's answer
2020-03-06T10:45:08-0500

Area of square is "A(s)=s^2."

To determine the rate change of its area with respect s, take the derivative of A.

"\frac{dA}{ds}=2s."

Note it is the same as A'. So,

"A'(s)=2s."

Then, substitute the given value of s.

"A'(8)=2(8)=16."

Answer:

The rate of change of area with respect to "s" when "s=8" is 16.